OpenStudy (anonymous):

The coordinates below represent two linear equations. How many solutions does this system of equations have? Line 1 Line 2 x y x y -3 6 -6 7 3 4 9 2 A. 0 B. exactly 1 C. exactly 2 D. infinitely many

OpenStudy (tkhunny):

Have you considered calculating the respective slopes?

OpenStudy (anonymous):

I have both of their slopes are \[\frac{ 1 }{ 3 }\]

OpenStudy (tkhunny):

Okay, that narrows it down to Zero (0) or Infinitely many. One (1) is no longer an option. Now what?

OpenStudy (tkhunny):

BTW c. Exactly 2 is just silly. Never pick that for linear equations.

OpenStudy (tkhunny):

You may also wish to figure out why you managed a positive slope when both are negative.

OpenStudy (anonymous):

\[\frac{ y _{1}-y _{2} }{ x _{1}-x _{2} }\]

OpenStudy (tkhunny):

Indeed. Somehow, something went wrong in there.

OpenStudy (anonymous):

Line 1 \[\frac{ 6-4 }{ -3-3 }\]

OpenStudy (tkhunny):

\(\dfrac{6-4}{-3-3} = \dfrac{2}{-6} = -\dfrac{1}{3}\)

OpenStudy (anonymous):

Oh right I forgot to include the - but it was in my notebook

OpenStudy (tkhunny):

Okay, now that we have the right equal slopes, we'll need the y-intercepts, as well. Are they the same line? Infinitely many solutions. Are tehy different lines? Zero solutions.

OpenStudy (anonymous):

I think the slope fr line 1 is 4 and the slope for line 2 is 2

OpenStudy (anonymous):

Is that right? I'm not sure.

OpenStudy (tkhunny):

?? We already established that the slopes of both are -1/3. What are these "4" and "2"?

OpenStudy (tkhunny):

Using the Point-Slope Form: y-6 = (-1/3)(x+3) y-7 = (-1/3)(x+6) Same or different?

OpenStudy (anonymous):

I meant the y-intercept

OpenStudy (anonymous):

They are different

OpenStudy (tkhunny):

y-6 = (-1/3)(x+3) y-7 = (-1/3)(x+6) -3y+18 = x+3 -3y+21 = x+6 15 = x + 3y 15 = x + 3y They look kind of the same to me. Why did you get a different result? You weren't being sloppy, were you?

OpenStudy (anonymous):

Oh, I messed up

OpenStudy (anonymous):

Thank you, you helped a ton!